5. A boy weighing 60kg slides on an ice rink at a speed of 4m / s. How fast will the boy slide if he throws a stone weighing 2.5kg horizontally behind him, at a speed of 1.5m / s. Friction is not taken into account.


1
Expert's answer
2021-02-08T18:42:35-0500
"Mv=MV-mu\\(60)(4)=(60)V-(2.5)(1.5)\\V=4.0625\frac{m}{s}"